If a ball is projected vertically upward then the distance travelled by the ball in last second ?

Grade:10

3 Answers

deekshit
8 years ago
let us consider a ball moves height h in t-1 seconds 
then h=u(t-1)-1/2g(t-1)^2.
if max height reached is H in t  seconds.
then H=ut-1/2gt^2.
subracting both u’ll get H-h=u+g/2(1-2t).
 
deekshit
8 years ago
substituting t=1 we will get correct answer i.e.u-g/2.
this gives correct answer.substituting t=1 is misssed in my answer.i want to known why my answer is treated as disapproved as i gave clear explanation
 
 
 
 
 
 
 
deekshit
8 years ago
ANOTHER METHOD DSTANCE TRAELLED IN NTH SECOND IS( S=U+a(n-1/2))
from this also we can get u-g/2. as u=u,a=-g,n=1.here we take n=1 because 
distance travelled in  first second of upward journey = distance travelled in last second of journey.